OpenStudy (anonymous):

Complete the square to solve for x. x^2+4x−21=0

OpenStudy (tkhunny):

Well, do it! (x^2 + 4x + ______) - 21 - _______ = 0 The blanks are the same.

OpenStudy (anonymous):

x^2+4x=21 x^2+4x+___=21+___ x^2+4x+4=21+4 (x+2)^2=25 x+2=+ or - sqrt25 x=-2 + sqrt 25 or x = -2 - sqrt 25 x=-2+5 or x -2-5 x=3 or x = -7

OpenStudy (anonymous):

thank you, i got the same thing!